OpenStudy (anonymous):

x - y = 4 x + y = 8 The x-coordinate of the solution to the system shown is _____. 2 4 6 12

OpenStudy (anonymous):

Solve x−y=4;x+y=8 Steps: I will try to solve your system of equations. x−y=4;x+y=8 Step: Solve x−y=4 for x: x−y+y=4+y(Add y to both sides) x=y+4 Step: Substitute (y+4) for x in x+y=8: x+y=8 y+4+y=8 2y+4=8(Simplify both sides of the equation) 2y+4+−4=8+−4(Add -4 to both sides) 2y=4 2y2=42(Divide both sides by 2) y=2 Step: Substitute (2) for y in x=y+4: x=y+4 x=2+4 x=6(Simplify both sides of the equation) Answer: x=6 and y=2